Sanjay Bhatnagar, Scientific Officer (SD), NCRA-TIFR, Pune University Campus, Pune 411 007.

July 13, 2001

The Head, Department of Physics, University of Pune, Ganeshkhind, Pune 411 007.

Dear Sir,

Enclosed herewith are forty five copies of the abstract of my Ph.D. thesis entitled “Radio Study of Galactic Supernova Remnants and the Interstellar Medium”. The research work presented in the thesis was carried out at the National Center for Radio Astrophysics (NCRA), Pune, under the guidance of Prof.D.J. Saikia, during the period of 1998 to 2001.

Thanking you.

Yours Sincerely

Sanjay Bhatnagar,

(Ph.D. Reg. No.: 88/98 /17063-65)